Pressure is often measured in units of atmosphere. To what depth must you swim to experience 3.0 atm pressure?

1
Expert's answer
2022-01-06T09:07:03-0500
"P=\rho gh,""h=\dfrac{P}{\rho g}=\dfrac{3\ atm\times\dfrac{101325\ Pa}{1\ atm}}{1000\ \dfrac{kg}{m^3}\times9.8\ \dfrac{m}{s^2}}=31\ m."